# How Can Time Series Analysis Benefit From Multiple Modalities? A Survey and Outlook

### v2

@@ 5.5. Domain-Specific Time-Series Works
MedTime 2 Text. This type of cross-modality work involves converting medical time-series data (e.g., ECG, ICU monitoring signals) into structured natural language reports for diagnostic and clinical interpretation. Specific application scenarios include: (1) ECG Report Generation generating cardiology reports from ECG time-series data (Bleich et al., 2024; Wan et al., 2024b; Tang et al., 2024b; Wan et al., 2024a; Li et al., 2024a) (2) ICU/NICU Data Summarization converting multichannel ICU/NICU physiological time-series data into summaries (Hunter et al., 2008b; Gatt et al., 2009; Hunter et al., 2008a) (3) Discharge Summary Automation generating coherent discharge summaries from time-series hospital course records (Hartman et al., 2023). Text 2 MedTime. In the medical domain, due to privacy concerns and limited medical resources, existing work focuses on using clinical text reports to synthesize realistic physiological signals, especially ECG (Alcaraz and Strodthoff, 2023; Chung et al., 2023; Lai et al., 2025). MedTime 2 Text + Text 2 MedTime. This type of work focuses on building chat models for medical time series analysis. ECG-Chat (Zhao et al., 2024b), ECG-Chat (Guo et al., 2024) and ECG Semantic Integrator (Yu et al., 2024) focus on QA for ECG signals; BioSignal Copilot (Liu et al., 2023c) further covers QA for multiple biosignals including ECG, EEG, PPG, EMG. MedTsLLM (Chan et al., 2024) can perform semantic segmentation, boundary detection and anomaly detection in time series, and includes text prompts to incorporate patient-specific information for deeper analysis of various physiological signals. Additionally, Cosentino et al. builds medical models for personal health.
@@ 5.5.3. Spatial-Temporal Analysis
Spatial-Temporal Data 2 Text. Unlike general TSA tasks, there haven’t been progress around generating other modalities regarding spatial-temporal time-series input, or vice versa. The closest progress is around the language-based analysis of sequential images and events, including satellite image series (Irvin et al., 2024; Yan et al., 2024) and urban event trajectories (Li et al., 2024b). At the same time, recent advances in adapting LLMs to time-series data, as aforementioned in Section 4.1 and Section 4.3, is capable of handling input from other modalities or adapting models from other modalities, but still targets on scenarios with time-series in both input and outputs. Instead, we argue that a comprehensive generative model, capable of handling input (or output) with modalities except time-series, are necessary for the development of spatial-temporal tasks as they can bridge the two modalities better. Spatial-Temporal Data 2 Image. Considering the complexity of spatiotemporal data, some studies focus on visualization solutions (Wang et al., 2017; Yang et al., 2019).
@@ Appendix G Details of Spatial-Temporal As Image
Yao et al. (2018) propose to forecast taxi demand by combining temporal view, spatial view, and semantic view, where spatial view is encoded by convolutional neural networks (CNNs). Li et al. (2022) use 3d convolutional networks to encode traffic trajectories and urban spatial correlations efficiently. Bao et al. (2022) introduce generator-discriminator structure to predict human mobility responses in COVID. Jiang et al. (2023) advance traffic forecasting by jointly encoding inputs by both graph neural networks and vision convolutional networks. Tang et al. (2024a) combine vision Mamba with LSTM and produce a strong model capable of predicting spatial taxi demand. These methods are strong in encoding spatial information. Similar to LLM applications, they do not involve extra input information, but treat the original inputs in an alternative manner.
@@ Appendix H More Details of Early Fusion
Each modality combination serves specific medical tasks. Time-series and text integration supports tasks such as mortality prediction, hospital readmission forecasting, disease phenotyping, and intensive care unit (ICU) deterioration modeling, where unstructured clinical notes complement physiological data in patient risk assessment (Lee et al., 2023; Wang et al., 2024c). Time-series and imaging fusion is commonly used in radiology for disease diagnosis and longitudinal tracking, such as leveraging sequential patient vitals and chest X-rays for pneumonia detection, pulmonary condition monitoring, and cardiovascular disease assessment (Wang et al., 2024c; Yao et al., 2024; Hayat et al., 2022). Time-series and tabular data fusion facilitates structured health risk modeling, as seen in treatment response prediction, where static baseline patient features (e.g., genetics, demographics) interact with dynamic disease progression markers (Soenksen et al., 2022; Zhang et al., 2023a; Kyung et al., 2024).

### v3

@@ 5.5. Domain-Specific Time-Series Works
MedTime 2 Text. This type of cross-modality work involves converting medical time-series data (e.g., ECG, ICU monitoring signals) into structured natural language reports for diagnostic and clinical interpretation. Specific application scenarios include: (1) ECG Report Generation generating cardiology reports from ECG time-series data (Bleich et al., 2024; Wan et al., 2024b; Tang et al., 2024b; Wan et al., 2024a; Li et al., 2024b) (2) ICU/NICU Data Summarization converting multichannel ICU/NICU physiological time-series data into summaries (Hunter et al., 2008b; Gatt et al., 2009; Hunter et al., 2008a) (3) Discharge Summary Automation generating coherent discharge summaries from time-series hospital course records (Hartman et al., 2023). Text 2 MedTime. In the medical domain, due to privacy concerns and limited medical resources, existing work focuses on using clinical text reports to synthesize realistic physiological signals, especially ECG (Alcaraz and Strodthoff, 2023; Chung et al., 2023; Lai et al., 2025). MedTime 2 Text + Text 2 MedTime. This type of work focuses on building chat models for medical time series analysis. ECG-Chat (Zhao et al., 2024b), ECG-Chat (Guo et al., 2024) and ECG Semantic Integrator (Yu et al., 2024) focus on QA for ECG signals; BioSignal Copilot (Liu et al., 2023c) further covers QA for multiple biosignals including ECG, EEG, PPG, EMG. MedTsLLM (Chan et al., 2024) can perform semantic segmentation, boundary detection and anomaly detection in time series, and includes text prompts to incorporate patient-specific information for deeper analysis of various physiological signals. Additionally, Cosentino et al. builds medical models for personal health.
@@ 5.5.3. Spatial-Temporal Analysis
Spatial-Temporal Data 2 Text. Unlike general TSA tasks, there haven’t been progress around generating other modalities regarding spatial-temporal time-series input, or vice versa. The closest progress is around the language-based analysis of sequential images and events, including satellite image series (Irvin et al., 2024; Yan et al., 2024) and urban event trajectories (Li et al., 2024c). At the same time, recent advances in adapting LLMs to time-series data, as aforementioned in Section 4.1 and Section 4.3, is capable of handling input from other modalities or adapting models from other modalities, but still targets on scenarios with time-series in both input and outputs. Instead, we argue that a comprehensive generative model, capable of handling input (or output) with modalities except time-series, are necessary for the development of spatial-temporal tasks as they can bridge the two modalities better. Spatial-Temporal Data 2 Image. Considering the complexity of spatiotemporal data, some studies focus on visualization solutions (Wang et al., 2017; Yang et al., 2019).
@@ Appendix G Details of Spatial-Temporal As Image
Yao et al. (2018) propose to forecast taxi demand by combining temporal view, spatial view, and semantic view, where spatial view is encoded by convolutional neural networks (CNNs). Li et al. (2022) use 3d convolutional networks to encode traffic trajectories and urban spatial correlations efficiently. Bao et al. (2022) introduce generator-discriminator structure to predict human mobility responses in COVID. Jiang et al. (2023) advance traffic forecasting by jointly encoding inputs by both graph neural networks and vision convolutional networks. Tang et al. (2024a) combine vision Mamba with LSTM and produce a strong model capable of predicting spatial taxi demand. These methods are strong in encoding spatial information. Similar to LLM applications, they do not involve extra input information, but treat the original inputs in an alternative manner. Besides, neural signals  (Wang et al., 2024b, 2023; Li et al., 2024a) can also be regarded as spatial-temporal data.
@@ Appendix H More Details of Early Fusion
Each modality combination serves specific medical tasks. Time-series and text integration supports tasks such as mortality prediction, hospital readmission forecasting, disease phenotyping, and intensive care unit (ICU) deterioration modeling, where unstructured clinical notes complement physiological data in patient risk assessment (Lee et al., 2023; Wang et al., 2024d). Time-series and imaging fusion is commonly used in radiology for disease diagnosis and longitudinal tracking, such as leveraging sequential patient vitals and chest X-rays for pneumonia detection, pulmonary condition monitoring, and cardiovascular disease assessment (Wang et al., 2024d; Yao et al., 2024; Hayat et al., 2022). Time-series and tabular data fusion facilitates structured health risk modeling, as seen in treatment response prediction, where static baseline patient features (e.g., genetics, demographics) interact with dynamic disease progression markers (Soenksen et al., 2022; Zhang et al., 2023a; Kyung et al., 2024).